f(x)= ax^2+bx+5
Find a and b if f(x+1)-f(x)=8x+3
Can someone please help me with this problem ive never learned how to do a problem like this and i have to have it done for tomorrow. Your help would be greatly appreciated!

Answer by Alan3354(69443)

f(x)= ax^2+bx+5
Find a and b if f(x+1)-f(x)=8x+3
------------
Find f(x+1)
f(x+1) = a*(x+1)^2 + b*(x+1) + 5
= ax^2 + 2ax + a + bx+b +5
ax^2 + bx +5 Subtract f(x)
---------------------------
= 2ax + a + b
That's = 8x+3
2a = 8
a = 4
--------
a+b = 3
b = -1